OpenStudy (anonymous):

help please!!! Find the measures of two angles, one positive and one negative, that are coterminal with pi/2.

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

AWe dam i hate these ._.

OpenStudy (anonymous):

ik they're really confusing!!!!

OpenStudy (here_to_help15):

lol

undeadknight26 (undeadknight26):

This might be a lot higher than my highest intellect sorry ._.

OpenStudy (anonymous):

ok thanks anyway

OpenStudy (here_to_help15):

@undeadknight26

OpenStudy (here_to_help15):

I need some help in Earth Science mind helping me while i help @em2000

OpenStudy (rational):

simply add or subtract `2pi` to get a coterminal angle

OpenStudy (rational):

pi/2 + 2pi = ?

OpenStudy (here_to_help15):

to get co-terminal angles add 360 to that angle and subtract 360 from that angle

OpenStudy (here_to_help15):

or in radians, add and subtract 2pi

OpenStudy (here_to_help15):

so calculate pi/2 + 2pi =... pi/2 - 2pi =...

OpenStudy (here_to_help15):

looking at the choice, you don't need to actually calculate them!

OpenStudy (anonymous):

is it b?

OpenStudy (here_to_help15):

Answer choices?

OpenStudy (here_to_help15):

Post them real fast ill be back

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

they're above in an attachment

OpenStudy (here_to_help15):

0/0 ok

OpenStudy (here_to_help15):

there is one choice which says to add and subtract 360 degrees :)

OpenStudy (anonymous):

@Here_to_Help15 thank u!!

OpenStudy (here_to_help15):

Np : )